Hi I need your help. Three consecutive odd integers are such that the square of the third is 264 more than the square of the second. Find the three integers. I'm trying to understand how you would write this out to be able to work it. Thank you, Reuben
Answer by edjones(8007)

Let the integers be x, x+2, x+4
(x+4)^2=(x+2)^2+264
x^2+8x+16=x^2+4x+4+264
4x-252=0 subtract the right side from the left.
4x=252
x=63
63, 65, 67 are the integers.
.
Ed